OpenStudy (anonymous):

Help with inequalities? 10 is no more than 4 times the sum of twice a number and 3. Define variable, write inequality, and solve

OpenStudy (texaschic101):

10 <= 4(2x + 3) -- distribute through the parenthesis 10 <= 8x + 12 -- subtract 12 from both sides 10 - 12 <= 8x -2 <= 8x -- divide by 8 -2/8 <= x -1/4 <= x I think .....